A. Determine if 1781 is divisible by 3, 6, 7, 8, and 9. (5 items x 2 points)


B. Determine if each of the following numbers is a prime or composite.


6. 828


7. 1666


8. 1781


9. 1125


10. 107


C. Find the greatest common divisor of each of the following pairs of integers.


11. 60 and 100


12. 45 and 33


13. 34 and 58


14. 77 and 128


15. 98 and 273


D. Find the least common multiple of each of the following pairs of integers.


16. 72 and 108


17. 175 and 245


18. 150 and 70


19. 32 and 27


20. 540 and 504


1
Expert's answer
2022-06-07T15:04:59-0400

A. The number 1781 is odd. Then it is not divisible by 6 and 8.

Determine if 1781 is divisible by 3, 6, 7, 8, and 9.

The sum of digits of the number is "1+7+8+1=17." 17 is not divisible by 3. Then 1781 is not divisible by 3, 6, and 9.


"1781=200(7)+50(7)+4(7)+3"

1781 is not divisible by 7.

1781 is not divisible by 3, 6, 7, 8, and 9.


B.

6. 828 is even => is divisible by 2 => is a composite.


7. 1666 is even => is divisible by 2 => is a composite.


8. 1781=13(137) => is divisible by 13 => is a composite.


9. 1125 is divisible by 5 => is a composite.


10. 107 is the 28th prime number.


"10^2<107<11^2"

107 is not divisible by 2, 3, 4, 5, 6, 7, 8,9, 10, and 11=> is a prime.


C. Find the greatest common divisor of each of the following pairs of integers.


11.

"60=2(2)(3)(5)"


"100=2(2)(5)(5)"

The greatest common divisor "2(2)(5)=20"


12.

"45=3(3)(5)"


"33=3(11)"

The greatest common divisor "3"


13.

"34=2(17)"


"58=2(29)"

The greatest common divisor "2"


14.

"77=7(11)"


"128=2(2)(2)(2)(2)(2)(2)"

The greatest common divisor "1"


15.

"98=2(7)(7)"




"273=3(7)(13)"

The greatest common divisor "7"


D. Find the least common multiple of each of the following pairs of integers.

16.

"72=2(2)(2)(3)(3)"




"108=2(2)(3)(3)(3)"

The least common multiple "2(2)(2)(3)(3)(3)=216"


17.

"175=5(5)(7)"




"245=5(7)(7)"

The least common multiple "5(5)(7)(7)=1225"


18.

"150=2(3)(5)(5)"




"70=2(5)(7)"

The least common multiple "2(3)(5)(5)(7)=1050"


19.

"32=2(2)(2)(2)(2)"




"27=3(3)(3)"

The least common multiple "2(2)(2)(2)(2)(3)(3)(3)=864"


20.

"540=2(2)(3)(3)(3)(5)"




"504=2(2)(2)(3)(3)(7)"

The least common multiple "2(2)(2)(3)(3)(3)(5)(7)=7560"